How do I solve
8/9z+2=1/3(4z-6)

Answer by stanbon(75887)

-----
8/9z+2=1/3(4z-6)
Multiply thru by 9 to get:
8z + 18 = 12z-18
4z = 36
z = 9
=======
Cheers,
Stan H.
==========